M&Ms are blended in a ratio of 9 percent brown, 13 percent yellow, 10 percent red, 20 percent blue, 17 percent orange, and 31 percent green. Suppose you choose a sample of two M&Ms at random from a large bag. (a) How many elements are there in the sample space, assuming the order in which the M&Ms are drawn matters? Sample space (b) What is the probability that both are brown? (Round your answer to 4 decimal places.) Probability .0169 (c) What is the probability that both are blue? (Round your answer to 4 decimal places.) Probability (d) What is the probability that both are green? (Round your answer to 4 decimal places.) Probability (e) What is the probability that the first M&M is brown and the second is green? (Round your answer to 4 decimal places.) Probability
Answer #1

total colors of M&Ms = 6

a) total elements in sample space = 6² = 36

b) P(both brown) = 0.09*0.09 = 0.0081

c) P(both blue) = 0.20*0.20=   0.04
d) P(both green) = 0.31*0.31=   0.0961

e) P(first brown and second green) = 0.09*0.31= 0.0279